### v2

@@ Abstract
In this paper, we study the significance of ecological interactions and separation of birth and death dynamics in heterogeneous stochastic populations via general birth-death processes. Interactions can manifest through the birth dynamics, the death dynamics, or some combination of the two. The underlying mechanisms are important but often implicit in population-level data. We propose an inference method for disambiguating the types of interaction and the birth and death processes from population size time series data of a stochastic $n$-type heterogeneous population. The interspecies interactions considered can be competitive, antagonistic, or mutualistic. We show that different pairs of birth and death rates with the same net growth rate result in different time series statistics. Then, the inference method is validated in the example of a birth-death process inspired by the two-type Lotka-Volterra interaction dynamics. Utilizing stochastic fluctuations enables us to estimate additional parameters in this stochastic Lotka-Volterra model, which are not identifiable in a deterministic model.
@@ 1 Introduction
This work focuses on understanding how the individual-level interactions manifest in the birth versus the death dynamics. Extracting this information is important in various applications, such as disambiguating exploitation versus interference competition in ecology, -static versus -cidal drug effects, evolution, and the hallmarks of cancers [1, 2, 3, 4, 5, 6].
In this paper, we build on the work of Huynh et al. 2023 [1], which focused on understanding birth versus death effects for homogeneous populations. Here, we extend the framework to heterogeneous populations with a particular emphasis on their ecological interactions. Our approach is inspired by the work of Cho et al. 2023 [7], which highlights the importance of different types of heterogeneous ecological interactions such as competitive, coooperative, and antagonistic. Interactions in heterogeneous populations have significant biological impact in various mechanisms including cancer therapy resistance [8, 9, 10], antibacterial resistance [11, 12, 13], and ecological adaptations [14] such as bet-hedging [15] and metapopulation migration [16].
The underlying model for the inference method is a general birth-death process. A birth-death process is a continuous-time Markov chain tracking the subpopulation sizes over time. The model is non-spatial and interactions between individuals are incorporated by allowing the birth and death rates to be dependent upon the population size in a nonlinear way.
@@
Other work has addressed the specific question of birth and death disambiguation, such as [17, 18, 19]. Crawford et al. 2014 [17] introduce an expectation-maximization algorithm to infer parameters in a general homogeneous birth-death process when the population size is sampled at discrete times. Their method involves a likelihood, which requires explicitly the form of the birth and the death rates. Here, we also assume that subpopulation sizes are sampled at discrete times, but unlike [17], we require that the time between samples be sufficiently small. The method we present applies even when the functional form of the birth and death rates is initially unknown. Roney et al. 2020 [18] use the maximum likelihood approach to estimate birth and death rates in a multi-type branching process and demonstrate their method in the case of linear birth and death rates. Their model does not include interaction between different populations. Gunnarsson et al. 2023 [19] also study multi-type branching processes with a focus on inferring phenotype switching between the subpopulations. Unlike these prior studies, our method applies to general, nonlinear birth and death rates, which allows interactions between populations through these rates.
@@ 2.1 General Birth-Death Process
Let ${\bm{e}}_{k}$ be the $k$th standard basis vector. For states $\bm{i},\bm{j}\in\mathbb{Z}^{n}$, the process has the infinitesimal transition probabilities given by
@@
for $\Delta t$ sufficiently small. We require that $\lambda_{k},\mu_{k}\geq 0$ and that $\lambda_{k}$ and $\mu_{k}$ are both 00 when $N_{k}=0$. We also assume that the rates are such that the process does not explode in finite time (that is, that there are not an infinite number of transitions in a finite time interval). However, we do not require these rates to be linear, and the rate at which type $k$ gives birth or dies can depend on the number of individuals of any subtype. In this way, the subpopulations interact with each other through the birth and the death rates. Both intraspecies interactions and interspecies interactions are possible. Figure 1 depicts a graphical representation when $n=2$. See [20] for a reference developing and discussing these types of processes in the context of biological applications or [21] for a more general introduction to continuous-time Markov chains.
@@ 2.2 Two-Type Lotka-Volterra Birth-Death Process
Throughout the paper, we will focus on a particular example of a birth-death process that falls within the general framework where the birth and death rates are inspired by the Lotka-Volterra system of differential equations.
@@ 3 Data Description
We use the tau-leaping algorithm to simulate the population time series. In this algorithm, the birth-death process is approximated by updating the birth and death rates at deterministic intervals of size $\tau>0$, rather than after each birth and death event. With this approximation, all events of a particular type occurring in the $\tau$-interval are simulated together using a Poisson random variable, rather than each event being tracked individually. This is an established stochastic simulation algorithm which speeds up computation of the time series while still maintaining reasonable accuracy. The error associated with the tau-leaping method is of order $\tau$; see [22] for an analysis of this method and comparison to other simulation algorithms. In generating the data for analysis, we used $\tau=0.1$. Using the tau-leaping scheme has the added benefit of creating data which is of the same form as datasets typically collected in laboratories [23], [24], [25], [26] or clinics. In these settings, it can be difficult to track individual cells in experiments or patients and instead, population counts are updated at fixed time intervals.
@@ 4.1 Time Series Statistics
Figure 2: In-silico time series of sensitive (blue) and resistant (red) populations of PC3 (a) and DU145 (b) cell lines for different $\gamma_{R}$ and $\sigma_{R}$ values. The variance is most affected by $\gamma_{R}$, i.e., the regulation of intraspecies competition on either birth or death.
@@
In case of PC3 in Figure 2(a), the capacity of sensitive cells is the same regardless of whether or not the resistant cells survive because the interaction parameter $\alpha_{S}$ is close to being trivial, $\alpha_{S}\approx 0$. With the DU145 parameters, we see in Figure 2(b) that the sensitive cells of DU145 either stay at the sensitive cell carrying capacity, corresponding to the runs when the resistant cells go extinct, or they reach a higher level when the resistant cells survive. This increased carrying capacity is due to the interaction type where the resistant population boosts the sensitive cells, $\alpha_{S}<0$. The third panel for the DU145 case shows a phenomenon that if the resistant population survives, then the number of sensitive cells eventually becomes constant. This is because at this point, the birth rate and death rate of the sensitive cells are both zero. In particular, though the choice of $\sigma_{k}$ and $\gamma_{k}$ do not affect the net growth rate of the subpopulations for most parameter values, if the maximum in Equations (13)-(16) forces a birth or death rate to be zero, then the net growth rate can be impacted by the choice of $\gamma_{k}$ or $\sigma_{k}$. In this case, the net growth rate becomes zero in a range where other choices of $\gamma_{k}$ and $\sigma_{k}$ would have a nonzero net growth rate.
@@
(a) PC3 : $\mathbf{\alpha_{S}=0.027,\alpha_{R}=0.159}$
(b) DU145 : $\mathbf{\alpha_{S}=-0.501,\alpha_{R}=0.221}$
Figure 3: Box plot statistics of sensitive (left) and resistant (right) populations of PC3 (a) and DU145 (b) cell lines at the final simulation time $t=100$, given that the resistant cells survive. The range of the number of cells at the final time is larger when $\gamma_{R}=0$ compared to $\gamma_{R}=1$, that is, when the intraspecies competition regulates death rather than birth. In the case of DU145, the median of the population differs from the other cases when $\gamma_{R}=1$ and $\sigma_{R}=0$.
@@ 4.2 Computational Survival Probability
In this section, we study the effect of birth- versus death-regulated interactions on the survival of the resistant population for different interspecies interaction regimes. We compute the probability that the resistant population does not go extinct for a fixed time period of $T=100$ when the population begins with one resistant cell. We refer to this the survival probability. This quantity demonstrates how likely the resistant population is to grow to carrying capacity, given that it begins with a single cell. The survival probability is numerically computed from $\mathbf{M=}$ 10000 time series by computing the proportion of runs in which the resistant population survived until time $T=100$. The simulation is initiated with $[S_{0},R_{0}]=[K_{S}-1,1]$ as in Section 4.1. The proportion of the $M$ runs with survival at time $T$ is the approximate survival probability:
$\mathbf{P\left[R(T)\neq 0|S(0)=K_{S}-1,R(0)=1\right]\approx\frac{1}{M}\sum_{i=1}^{M}\mathbbm{1}_{R^{i}(T)\neq 0},}$ (17)
@@
(a) PC3 ($\mathbf{\alpha_{S}=0.027,\alpha_{R}=0.159}$)                     (b) DU145 ($\mathbf{\alpha_{S}=-0.501,\alpha_{R}=0.221}$)
@@
(c) case 3 ($\mathbf{\alpha_{S}=0.159,\alpha_{R}=0.027}$)                       (d) case 4 ($\mathbf{\alpha_{S}=0.221,\alpha_{R}=-0.501}$)
@@
Figure 4 plots the survival probability of the resistant population for different values of $\gamma_{R}=0,0.1,...,1$ and $\sigma_{R}=0,0.1,...,1$. In the cases of PC3 and DU145, the survival probability is more sensitive to $\sigma_{R}$ compared to $\gamma_{R}$. The birth versus death regulation of interspecies interaction affects the survival probability more than the intraspecies competition. The survival probability in both PC3 and DU145 increases about 7% as $\sigma_{R}$ increases from 0 to 1. Both cases are when $\alpha_{R}>0$. In addition, we test two different types of interaction where $\alpha_{R}$ has different signs, $\mathbf{\alpha_{R}\approx 0}$ in Figure 4(c) and $\mathbf{\alpha_{R}<0}$ in Figure 4(d). When $\mathbf{\alpha_{R}\approx 0}$, the effect of the sensitive cells on the resistant cells is small, thus $\sigma_{R}$ is not applicable and the survival probability stays relatively consistent regardless of $\sigma_{R}$. However, when $\alpha_{R}<0$ in Figure 4(d), the survival probability decreases as $\sigma_{R}$ increases. Thus, the sign of $\alpha_{R}$ changes the sign of the correlation between the survival probability and $\sigma_{R}$.
@@
The change in the survival probability of the resistant population across different values of $\sigma_{R}$ depends on the sign of $\alpha_{R}$ but not on the sign of $\alpha_{S}$. It is reasonable that the parameter that describes how the sensitive cells impact the resistant cells affects the survival probability of the resistant population more than the type of interaction in the opposite direction. When $\alpha_{R}=0.5$ and the resistant cells compete with the sensitives cells, the survival probability is around $0.55$ to $0.7$. However, when $\alpha_{R}=-0.5$ and the sensitives cells positively impact the resistant cells, the survival probability increases to the level of $0.75$ to $0.95$. In addition to the numerical probability of survival, we again confirm that the correlation of $\sigma_{R}$ to the survival probability changes depending on the sign of $\alpha_{R}$. When $\alpha_{R}>0$, the survival probability increases as $\sigma_{R}$ increases, but when $\alpha_{R}<0$, the survival probability decreases as $\sigma_{R}$ increases.
@@ 5.2 Birth and Death Rate Inference Algorithm
    First, we divide the space $\mathbb{R}^{\geq 0}\times\mathbb{R}^{\geq 0}$ into a grid. The grid can have a fixed grid size $\Delta x$ or can be chosen to vary to more appropriately capture the data.
@@
The choice of grid size is important. For the method to be accurate, there need to be enough data points in each grid block to get a good approximation of the mean and variance, so the grid blocks must be big enough to have a large number of points in each nonempty grid block. However, because $\lambda_{k},\mu_{k}$ are functions of ${\bm{N}}$, choosing the grid too large may group points together which have quite different rates. This will also reduce the accuracy of the method. In the homogeneous population case, the analysis of Huynh et al. 2023 [1] suggests that intermediate grid sizes are optimal. For the two-population Lotka-Volterra model, we analyze the error as a function of grid size $\Delta x$ in Figure 9.
@@ 5.4 Numerical Validation
Figure 6: Total birth rate $\lambda_{k}^{i,j}$ and death rate $\mu_{k}^{i,j}$ (dots) estimated by the method in Section 5.2, and the fitted birth and death rate functions, $b^{mono}_{k}(N_{k};\theta)$ and $d^{mono}_{k}(N_{k};\theta)$ for monoculture (a) and $b^{co}_{k}(N_{1},N_{2};\theta)$ and $d^{co}_{k}(N_{1},N_{2};\theta)$ for coculture (b), for $k=S,R$, determined by the method in Section 5.3. The estimated rate functions $b_{k}$ and $d_{k}$ exhibit the correct linear/nonlinear behavior. When $\gamma_{S}=\gamma_{R}=0,\,\sigma_{S}=\sigma_{R}=0$ (red), the birth rates are linear functions of $S$ and $R$, while death rates are nonlinear functions of $S$ and $R$. On the other hand, when $\gamma_{S}=\gamma_{R}=1,\,\sigma_{S}=\sigma_{R}=1$ (blue), the death rates are linear functions, and the birth rates are nonlinear functions.
In this section, we validate the proposed inference method on simulated data, comparing the inferred parameters with the true simulation parameters. In particular, we analyze the inference errors in terms of number of time series $M$, discretized time step $\Delta t$, and grid size $\Delta x$. Unless mentioned otherwise, the default values of the inference method parameters are chosen as follows: $M=100$, $\Delta x=10$, $\Delta t=0.1$. The bin exclusion threshold is set at $100$, where bins containing less than 100 data points are excluded. Since the subpopulation carrying capacities in our simulated data are on the order of $10^{3}$, the choice of grid size $\Delta x=10$ divides the time series into $O(10^{2})$ bins. The final time of the stochastic simulation is $T=100$; thus, $\Delta t=0.1$ yields 1000 discretized time points. For the other model parameters, we focus on the specific values associated with the cancer cell line PC3. Parameters of the PC3 cell line can be found in Table 2.
@@
Figure 7: Relative error of PC3 parameter estimation with respect to sample size $M=10,100,1000$. The error decreases as the sample size increases.
@@
Figure 8: Relative error of PC3 parameter estimation with respect to time step size $\Delta t=10,1,0.1$. The error decays as $\Delta t$ decreases.
Figure 8 shows the error with respect to $\Delta t$, the time step of stochastic simulation. The shown results are for $\Delta t=10,1$, and $0.1$, and the accuracy improves as $\Delta t$ decreases. Due to stochasticity in the time series, using a time step that is too small makes the rate inference less accurate. Moreover, it shows that using a large time step such as $\Delta t=10$ significantly deteriorates the accuracy.
Figure 9: Relative error of PC3 parameter estimation with respect to grid size $\Delta x=50,10,2$. The error is less sensitive to the grid size than other inference parameters.
In Figure 9, the grid sizes $\Delta x=50,10$, and $2$ are tested. The error seems to be the least sensitive to $\Delta x$ as compared to the other tested hyper-parameters $M$ and $\Delta t$.

### v3

@@ Abstract
In this paper, we study the significance of ecological interactions and separation of birth and death dynamics in stochastic heterogeneous populations via general birth-death processes. Interactions can manifest through the birth dynamics, the death dynamics, or some combination of the two. The underlying microscopic mechanisms are important but often implicit in population-level data. We propose an inference method for disambiguating the types of interaction and the birth and death processes from population size time series data of a stochastic $n$-type heterogeneous population. The interspecies interactions considered can be competitive, antagonistic, or mutualistic. We show that different pairs of birth and death rates with the same net growth rate result in different time series statistics. Then, the inference method is validated in the example of a birth-death process inspired by the two-type Lotka-Volterra interaction dynamics. Utilizing stochastic fluctuations enables us to estimate additional parameters in this stochastic Lotka-Volterra model, which are not identifiable in a deterministic model.
@@ 1 Introduction
This work focuses on understanding how individual-level interactions manifest in a process that involves birth and death dynamics, with motivation from ecological interactions. Extracting this information is important in many applications such as disambiguating exploitation versus interference competition in ecology, -static versus -cidal drug effects, evolution, and the hallmarks of cancers [1, 2, 3, 4, 5, 6].
In this paper, we build on the work of Huynh et al. 2023 [1], which focused on understanding birth versus death effects for homogeneous populations. Here, we extend the framework to heterogeneous populations with a particular emphasis on their ecological interactions. This extension is inspired by the work of Cho et al. 2023 [7], which highlights the importance of different types of heterogeneous ecological interactions such as competitive, cooperative, and antagonistic. Interactions in heterogeneous populations have significant biological impact in various mechanisms including cancer therapy resistance [8, 9, 10], antibacterial resistance [11, 12, 13], and ecological adaptations [14] such as bet-hedging [15] and metapopulation migration [16].
The underlying model for our inference method is a general birth-death process. A birth-death process is a continuous-time Markov chain tracking the subpopulation sizes over time. The model is non-spatial and interactions between individuals are incorporated by allowing the birth and death rates to be dependent upon the population size in a nonlinear way.
@@
Other work has addressed the specific question of birth and death disambiguation, such as [17, 18, 19]. Crawford et al. 2014 [17] introduce an expectation-maximization algorithm to infer parameters in a general homogeneous birth-death process when the population size is sampled at discrete times. Their method involves a likelihood, which requires explicitly the form of the birth and the death rates. Here, we also assume that subpopulation sizes are sampled at discrete times, but unlike [17], we require that the time between samples are sufficiently small. The method we present applies even when the functional form of the birth and death rates is initially unknown. Roney et al. 2020 [18] use the maximum likelihood approach to estimate birth and death rates in a multi-type branching process and demonstrate their method in the case of linear birth and death rates. Their model does not include interaction between different populations. Gunnarsson et al. 2023 [19] also study multi-type branching processes with a focus on inferring phenotype switching between the subpopulations. Unlike these prior studies, our method applies to general, nonlinear birth and death rates, which allows interactions between populations through these rates.
@@ 2.1 General Birth-Death Process
Let ${\bm{e}}_{k}$ be the $k$-th standard basis vector. For states $\bm{i},\bm{j}\in\mathbb{Z}^{n}$, the process has the infinitesimal transition probabilities given by
@@
for $\Delta t$ sufficiently small. We require that $\lambda_{k},\mu_{k}\geq 0$ and that $\mu_{k}=0$ when $N_{k}=0$. We also assume that the rates are such that the process does not explode in finite time (that is, that there are not an infinite number of transitions in a finite time interval). However, we do not require these rates to be linear, and the rate at which type $k$ gives birth or dies can depend on the number of individuals of any subtype. In this way, the subpopulations interact with each other through the birth and the death rates. Both intraspecies interactions and interspecies interactions are possible. Figure 1 depicts a graphical representation when $n=2$. See [20] for a reference developing and discussing these types of processes in the context of biological applications or [21] for a more general introduction to continuous-time Markov chains.
@@ 2.2 Two-Type Lotka-Volterra Birth-Death Process
Throughout the paper, we will focus on a particular example of a birth-death process that falls within this general framework where the birth and death rates are inspired by the Lotka-Volterra system of differential equations.
@@ 3 Data Description
We use the tau-leaping algorithm to simulate the population time series. In this algorithm, the birth-death process is approximated by updating the birth and death rates at deterministic intervals of size $\tau>0$, rather than after each birth and death event. With this approximation, all events of a particular type occurring in the $\tau$-interval are simulated together using a Poisson random variable, rather than each event being tracked individually. This is an established stochastic simulation algorithm which speeds up computation of the time series while still maintaining reasonable accuracy. The error associated with the tau-leaping method is of order $\tau$; see [22] for an analysis of this method and comparison to other simulation algorithms. In generating the data for analysis, we used $\tau=0.1$. Using the tau-leaping scheme has the added benefit of creating data which is of the same form as datasets typically collected in laboratories [23, 24, 25, 26] or clinics. In these settings, it can be difficult to track individual cells in experiments or patients and instead, population counts are updated at fixed time intervals.
@@ 4.1 Time Series Statistics
Figure 2: In-silico time series of sensitive (blue) and resistant (red) populations of (a) PC3 and (b) DU145 cell lines for different $\gamma_{R}$ and $\sigma_{R}$ values. The variance is most affected by $\gamma_{R}$, i.e., the regulation of intraspecies competition on either birth or death.
@@
In the case of PC3 in Figure 2(a), the capacity of sensitive cells is the same regardless of whether or not the resistant cells survive because the interaction parameter $\alpha_{S}$ is close to being trivial, $\alpha_{S}\approx 0$. With the DU145 parameters, we see in Figure 2(b) that the sensitive cells of DU145 either stay at the sensitive cell carrying capacity, corresponding to the runs when the resistant cells go extinct, or they reach a higher level when the resistant cells survive. This increased carrying capacity is due to the interaction type where the resistant population boosts the sensitive cells, $\alpha_{S}<0$. The third panel for the DU145 case shows a phenomenon that if the resistant population survives, then the number of sensitive cells eventually becomes constant. This is because at this point, the birth rate and death rate of the sensitive cells are both zero. In particular, though the choice of $\sigma_{k}$ and $\gamma_{k}$ does not affect the net growth rate of the subpopulations for most parameter values, if the maximum in Equations (13)-(16) forces a birth or death rate to be zero, then the net growth rate can be impacted by the choice of $\gamma_{k}$ or $\sigma_{k}$. In this case, the net growth rate becomes zero in a range where other choices of $\gamma_{k}$ and $\sigma_{k}$ would have a nonzero net growth rate.
@@
(a) PC3 : $\alpha_{S}=0.027,\alpha_{R}=0.159$
(b) DU145 : $\alpha_{S}=-0.501,\alpha_{R}=0.221$
Figure 3: Box plot statistics of sensitive (left) and resistant (right) populations of (a) PC3 and (b) DU145 cell lines at the final simulation time $t=100$, given that the resistant cells survive. The range of the number of cells at the final time is larger when $\gamma_{R}=0$ compared to $\gamma_{R}=1$, that is, when the intraspecies competition regulates death rather than birth. In the case of DU145, the median of the population differs from the other cases when $\gamma_{R}=1$ and $\sigma_{R}=0$.
@@ 4.2 Computational Survival Probability
In this section, we study the effect of birth- versus death-regulated interactions on the survival of the resistant population for different interspecies interaction regimes. We compute the probability that the resistant population does not go extinct for a fixed time period of $T=100$ when the population begins with one resistant cell. We refer to this as the survival probability. This quantity demonstrates how likely the resistant population is to grow to carrying capacity, given that it begins with a single cell. The survival probability is numerically computed from ${M=}$ 10000 time series by computing the proportion of runs in which the resistant population survived until time $T=100$. The simulation is initiated with $[S_{0},R_{0}]=[K_{S}-1,1]$ as in Section 4.1. The proportion of the $M$ runs with survival at time $T$ is the approximate survival probability:
$P\left[R(T)\neq 0|S(0)=K_{S}-1,R(0)=1\right]\approx\frac{1}{M}\sum_{i=1}^{M}\mathbbm{1}_{R^{i}(T)\neq 0},$ (17)
@@
(a) PC3 ($\alpha_{S}=0.027,\alpha_{R}=0.159$)                     (b) DU145 ($\alpha_{S}=-0.501,\alpha_{R}=0.221$)
@@
(c) case 3 ($\alpha_{S}=0.159,\alpha_{R}=0.027$)                       (d) case 4 ($\alpha_{S}=0.221,\alpha_{R}=-0.501$)
@@
Figure 4 plots the survival probability of the resistant population for different values of $\gamma_{R}=0,0.1,...,1$ and $\sigma_{R}=0,0.1,...,1$. In the cases of PC3 and DU145, the survival probability is more sensitive to $\sigma_{R}$ compared to $\gamma_{R}$. The birth versus death regulation of interspecies interaction affects the survival probability more than the intraspecies competition. The survival probability in both PC3 and DU145 increases about 7% as $\sigma_{R}$ increases from 0 to 1. Both cases are when $\alpha_{R}>0$. In addition, we test two different types of interaction where $\alpha_{R}$ has different signs, $\alpha_{R}\approx 0$ in Figure 4(c) and $\alpha_{R}<0$ in Figure 4(d). When $\alpha_{R}\approx 0$ the effect of the sensitive cells on the resistant cells is small, thus $\sigma_{R}$ is not applicable and the survival probability stays relatively consistent regardless of $\sigma_{R}$. However, when $\alpha_{R}<0$ in Figure 4(d), the survival probability decreases as $\sigma_{R}$ increases. Thus, the sign of $\alpha_{R}$ changes the sign of the correlation between the survival probability and $\sigma_{R}$.
@@
The change in the survival probability of the resistant population across different values of $\sigma_{R}$ depends on the sign of $\alpha_{R}$ but not on the sign of $\alpha_{S}$. It is reasonable that the parameter that describes how the sensitive cells impact the resistant cells affects the survival probability of the resistant population more than the type of interaction in the opposite direction. When $\alpha_{R}=0.5$ and the resistant cells compete with the sensitive cells, the survival probability is around $0.55$ to $0.7$. However, when $\alpha_{R}=-0.5$ and the sensitive cells positively impact the resistant cells, the survival probability increases to the level of $0.75$ to $0.95$. In addition to the numerical probability of survival, we again confirm that the correlation of $\sigma_{R}$ to the survival probability changes depending on the sign of $\alpha_{R}$. When $\alpha_{R}>0$, the survival probability increases as $\sigma_{R}$ increases, but when $\alpha_{R}<0$, the survival probability decreases as $\sigma_{R}$ increases.
@@ 5.2 Birth and Death Rate Inference Algorithm
    First, we divide the space $\mathbb{R}^{\geq 0}\times\mathbb{R}^{\geq 0}$ into a grid. The grid can have a fixed grid size $\Delta x$ or can be chosen to vary, to more appropriately capture the data.
@@
The choice of grid size is important. For the method to be accurate, there needs to be enough data points in each grid block to get a good approximation of the mean and variance, so the grid blocks must be big enough to have a large number of points in each nonempty grid block. However, because $\lambda_{k},\mu_{k}$ are functions of ${\bm{N}}$, choosing the grid too large may group points together which have quite different rates. This will also reduce the accuracy of the method. In the homogeneous population case, the analysis of Huynh et al. 2023 [1] suggests that intermediate grid sizes are optimal. For the two-population Lotka-Volterra model, we analyze the error as a function of grid size $\Delta x$ in Figure 9.
@@ 5.4 Numerical Validation
Figure 6: Total birth rate $\lambda_{k}^{i,j}$ and death rate $\mu_{k}^{i,j}$ (dots) estimated by the method in Section 5.2, and the fitted birth and death rate functions, $b^{mono}_{k}(N_{k};\theta)$ and $d^{mono}_{k}(N_{k};\theta)$ for (a) monoculture and $b^{co}_{k}(N_{1},N_{2};\theta)$ and $d^{co}_{k}(N_{1},N_{2};\theta)$ for (b) coculture, for $k=S,R$, determined by the method in Section 5.3. The estimated rate functions $b_{k}$ and $d_{k}$ exhibit the correct linear/nonlinear behavior. When $\gamma_{S}=\gamma_{R}=0,\,\sigma_{S}=\sigma_{R}=0$ (red), the birth rates are linear functions of $S$ and $R$, while death rates are nonlinear functions of $S$ and $R$. On the other hand, when $\gamma_{S}=\gamma_{R}=1,\,\sigma_{S}=\sigma_{R}=1$ (blue), the death rates are linear functions, and the birth rates are nonlinear functions.
In this section, we validate the proposed inference method on simulated data, comparing the inferred parameters with the true simulation parameters. In particular, we analyze the inference errors in terms of number of time series $M$, discretized time step $\Delta t$, and grid size $\Delta x$. Unless mentioned otherwise, the default values of the inference method parameters are chosen as follows: $M=100$, $\Delta x=10$, $\Delta t=0.1$. The bin exclusion threshold is set at $100$, where bins containing less than 100 data points are excluded. Since the subpopulation carrying capacities in our simulated data are on the order of $10^{3}$, the choice of grid size $\Delta x=10$ divides the time series into $O(10^{2})$ monoculture bins and $O(10^{4})$ coculture bins. The final time of the stochastic simulation is $T=100$; thus, $\Delta t=0.1$ yields 1000 discretized time points. For the other model parameters, we focus on the specific values associated with the cancer cell line PC3. Parameters of the PC3 cell line can be found in Table 2.
@@
Figure 7: Relative error of PC3 parameter estimation with respect to number of time series $M=10,100,1000$. The error decreases as the number of time series increases.
@@
Figure 8: Relative error of PC3 parameter estimation with respect to time step size $\Delta t=0.1,1,10$. The error decays as $\Delta t$ decreases.
Figure 8 shows the error with respect to $\Delta t$, the time step of stochastic simulation. The shown results are for $\Delta t=0.1,1$, and $10$, and the accuracy improves as $\Delta t$ decreases. Due to stochasticity in the time series, using a time step that is too small makes the rate inference less accurate. Moreover, it shows that using a large time step such as $\Delta t=10$ significantly deteriorates the accuracy.
Figure 9: Relative error of PC3 parameter estimation with respect to grid size $\Delta x=2,10,50$. The error is less sensitive to the grid size than other inference parameters.
In Figure 9, the grid sizes $\Delta x=2,10$, and $50$ are tested. The error seems to be the least sensitive to $\Delta x$ as compared to the other tested hyper-parameters $M$ and $\Delta t$. However, as mentioned above, the choice of grid size will influence which grid blocks have enough population points to compute the mean and variance; here, this cutoff is set at 100.